### v2

@@ 1 Introduction
In his classic essay, “The Normative Structure of Science,” sociologist Robert K. Merton identified universalism as a foundational principle of the scientific enterprise, one that distinguishes science from other competing systems of knowing. According to Merton and Storer’s formulation (Merton and Storer,, 1973, p. 270), universalism holds that the evaluation of scientific contributions “is not to depend on the personal or social attributes of their protagonist; his race, nationality, religion, class, and personal qualities are as such irrelevant.” The value of universalism is manifested perhaps most concretely in the practice of double-blind peer review, wherein the identities of both those making scientific claims and those evaluating them are obscured from one another (Bornmann,, 2011).
@@
While scholars have long observed that adherence to the principle of universalism is far from universal (Mulkay,, 1976; Cole,, 1992; Long and Fox,, 1995), the growing availability of large-scale databases is creating opportunities for unprecedented insight into processes of scientific evaluation (Teplitskiy et al.,, 2018; Dondio et al.,, 2019; Lane et al.,, 2021), including the barriers that inhibit objective assessments. Recent literature in particular has raised considerable concern about the role of gender in scientific evaluation (Moss-Racusin et al.,, 2012; Reuben et al.,, 2014; Oliveira et al.,, 2019; Card et al., 2020a, ). There is growing evidence to suggest that the contributions of women are devalued relative to men (Joshi,, 2014; Sarsons,, 2017; Vásárhelyi et al.,, 2021; King et al.,, 2017; Ni et al.,, 2021), thereby undermining the imperative of universalism, and ultimately weakening the scientific enterprise by suppressing diverse perspectives. Prior work has shown, for example, that relative to men, women’s contributions are awarded less grant funding from prestigious sponsors (Oliveira et al.,, 2019; Kolev et al.,, 2020), less likely to be granted patent protection (Jensen et al.,, 2018), and also receive fewer citations from other academic papers (Larivière et al.,, 2013). These gaps appear to be driven at least in part by homophily, as men tend to cite other men over women, even in fields with higher female representation, such as the social sciences (Potthoff and Zimmermann,, 2017; Dion et al.,, 2018; Ghiasi et al.,, 2018). Men also tend to self-cite at a much higher rate than women (King et al.,, 2017).
@@
Yet much remains unknown about the role of gender in scientific assessment. Prior scholarship has highlighted differences in field, topic choice, or research quality as playing an important role in explaining observed gaps between men and women (Leahey,, 2007; Key and Sumner,, 2019). While these factors are undoubtedly important, studies have also shown that even when accounting for field, topic choice, and quality, men’s and women’s research is evaluated differently (Hengel,, 2017; Hengel and Moon,, 2020). Studies further report gender gaps even in the presence of double-blind review (Kolev et al.,, 2020; Mahajan et al.,, 2020), suggesting that conscious or unconscious bias based on the identity of the author is likely not the determinative factor (van der Lee and Ellemers,, 2015; Hospido and Sanz,, 2019; Card et al.,, 2021; Card et al., 2020b, ).
@@
In this article, we aim to contribute to the systematic understanding of gender differences in the evaluation of scientific and technical work using a linguistic lens. Specifically, we draw on a prior framework, developed by Biber, (1988), which contrasts linguistic features characteristic of a high information density (“informational” features) with linguistic features associated with more affective, interactional content (“involved” features). This framework allows us to take a systematic approach to feature selection and also to examine features which have been overlooked as indicators of variation in the previous literature on gender gaps, such as the use of determiners or pronouns. Biber, (1988) found in a corpus of epistolary writing that women used more “involved” features (e.g. increased use of pronouns) for the purpose of building a relationship with the audience, while men used more “informational” features (e.g. increased use of cardinal numbers in texts) for the purpose of conveying the factual information. Using this framework, we conduct a large-scale analysis of gender differences in scientific and technical writing, leveraging two distinctive document types—academic articles and patented inventions—which collectively span all major fields of science and technology.
@@
Several noteworthy findings emerge from our study. We document that even in such restricted genres as academic writing and invention descriptions, female authors tend to use more “involved” features than male authors, a pattern that holds across all scientific fields and patent subcategories. In the patent data specifically, we find that the gender of the lawyer appears to have more impact on the writing than the gender of the inventor, suggesting both the importance of attorneys in crafting patent text and the universal nature of the patterns we observe, even across individuals with very different professional training. Finally, we show that work that uses a higher number of “involved” features is more likely to be cited by female authors.
@@
Our study builds on, but also extends, a small but growing stream of research in the Science of Science on gender and language (c.f., Fortunato et al.,, 2018). First, while existing literature has identified a number of differences in men’s and women’s scientific writing (Lerchenmueller et al.,, 2019; Kolev et al.,, 2020; Kim et al.,, 2022), the choice of features have not been motivated by a linguistic framework, thereby complicating the interpretation of the results and making comparisons across studies difficult. Second, existing research has been done within particular fields and on particular kinds of text (e.g., scientific papers), and therefore little is known about whether the patterns observed thus far will generalize across diverse fields and document types (such as patent data), which have diverse norms.
@@
Our study also adds to the linguistics literature. First, previous research on gender differences in language utilization has largely been done on informal or conversational text, typically focusing on speech patterns, and using corpora like emails (Thomson et al.,, 2001), essays (Mulac et al.,, 2006), social media (Garimella and Mihalcea,, 2016), blogs (Pennebaker,, 2011). There has been far less research on gender differences in formal writing, and the work that has been done has relied on relatively small samples (Argamon et al.,, 2003; Biber,, 1988). Second, we study a new source of text data—patented inventions—that has not been studied in previous linguistic analyses. Patents represent a more restricted genre in terms of stylistic expression than academic papers, therefore, finding stylistic differences by gender is especially notable.
@@ 2.1 Data Overview
We conduct the analysis on two datasets: abstracts of academic papers from Web of Science [^1] Web of Science: https://clarivate.com/webofsciencegroup/solutions/web-of-science/ and abstracts of patents from PatentsView [^2] PatentsView: http://www.patentsview.org/. For the purposes of our study, we narrow our focus to solo-authored papers and patents (i.e., we exclude papers and patents written by teams). In the patent data, we also narrow our sample to include only those inventors with a single lawyer, in order to more cleanly control for the gender of the lawyer, as lawyers are frequently involved in the writing and editing of patent text. In total, we use 256,260 publications and 16,053 patents in the analysis. Additional details on the dataset construction can be found in the Materials and Methods section.
For each paper and patent in our sample, we construct a series of measures tracking the utilization of informational features (using past tense, determiners, and cardinal numbers as indicators) and involved features (using pronouns, questions, and non-phrasal coordination (“and”) as indicators). Figure 1 gives an overview of the specific features of interest as well as more details about the computation of these features. [^3] We also released an open-source Python package to compute these linguistic features, available here: https://pypi.org/project/stylometer. We define three separate measures: involved rate, informational rate, and Involved-Informational Ratio. Because our corpus includes academic writing and invention descriptions, there is already a high utilization of informational features; therefore, in addition to involved and informational rates, we are also interested in seeing the relative rate at which writers use involved features in relation to informational features (Involved-Informational Ratio). More information on the developed measures is included in the Materials and Methods section.
@@ 2.2 Writing Examples
Although both abstracts address similar topics, they are presented in dramatically different ways. Abstract (2), written by the male author, is striking for its heavy utilization of cardinal numbers (e.g., “\$100 million”, “5.5 years”, “above 336.69”), which are mostly avoided by the female author. Abstract (1), by contrast, stands out for its greater utilization of “connector words” (e.g., “and”) and questions, the latter of which are not used at all in Abstract (2). The female abstract reads a lot like a story, building a relationship with the reader through the use of pronouns, present tense, asking questions, and by stringing the clauses together with “and” in a loose, storytelling-like manner; the male abstract is focused on relaying the factual information, via the use of numbers and determiners, and past tense. Figure 2 shows the full text of both abstracts, zoomed out, with informational and involved features highlighted in different colors (blue for informational, yellow for involved).

### v4

@@ 1 Introduction
In his classic essay, “The Normative Structure of Science,” sociologist Robert K. Merton identified universalism as a foundational principle of the scientific enterprise, one that distinguishes science from other competing systems of knowing. According to Merton’s formulation (merton1973, p. 270), universalism holds that the evaluation of scientific contributions “is not to depend on the personal or social attributes of their protagonist; his race, nationality, religion, class, and personal qualities are as such irrelevant.” The value of universalism is manifested perhaps most concretely in the practice of double-blind peer review, wherein the identities of both those making scientific claims and those evaluating them are obscured from one another (bornmann2011).
@@
While scholars have long observed that adherence to the principle of universalism is far from universal (mulkay1976; cole1992; longandfox1995), the growing availability of large-scale databases is creating opportunities for unprecedented insight into processes of scientific evaluation (teplitskiy2018; dondio2019; lane2021), including the barriers that inhibit objective assessments. Recent literature in particular has raised considerable concern about the role of gender in scientific evaluation (mossracusin2012; reuben2014; oliveira2019; card2020). There is growing evidence to suggest that the contributions of women are devalued relative to men (joshi2014; sarsons2017; vasarhelyi2021; king2017; ni2021), thereby undermining the imperative of universalism, and ultimately weakening the scientific enterprise by suppressing diverse perspectives. Prior work has shown, for example, that relative to men, women’s contributions are awarded less grant funding from prestigious sponsors (oliveira2019; kolev2020), less likely to be granted patent protection (jensen2018), and also receive fewer citations from other academic papers (lariviere2013). These gaps appear to be driven at least in part by homophily, as men tend to cite other men over women, even in fields with higher female representation, such as the social sciences (potthoff2017; dion2018; ghiasi2018). There is also some evidence that men tend to self-cite at a much higher rate than women (king2017), although findings are mixed and likely vary across fields (azoulay2020).
@@
Yet much remains unknown about the role of gender in scientific assessment. Prior scholarship has highlighted differences in field, topic choice, or research quality as playing an important role in explaining observed gaps between men and women (leahey2007; key2019). While these factors are undoubtedly important, studies have also shown that even when accounting for field, topic choice, and quality, men’s and women’s research is evaluated differently (hengel2017; hengel2020). Studies further report gender gaps even in the presence of double-blind review (kolev2020; mahajan2020), suggesting that conscious or unconscious bias based on the identity of the author is likely not the determinative factor (vanderlee2015; hospido2019; card2021; card2020_gender).
@@
In this article, we aim to contribute to the systematic understanding of gender differences in the evaluation of scientific and technical work using a linguistic lens. Specifically, we draw on a prior framework, developed by biber1988, which contrasts linguistic features characteristic of a high information density (“informational” features) with linguistic features associated with more affective, interactional content (“involved” features). This framework allows us to take a systematic approach to feature selection and also to examine features which have been overlooked as indicators of variation in the previous literature on gender gaps, such as the use of determiners or pronouns. biber1988 found in a corpus of epistolary writing that women used more “involved” features (e.g. increased use of pronouns) for the purpose of building a relationship with the audience, while men used more “informational” features (e.g. increased use of cardinal numbers in texts) for the purpose of conveying the factual information. Using this framework, we conduct a large-scale analysis of gender differences in scientific and technical writing, leveraging two distinctive document types—academic articles and patented inventions—which collectively span all major fields of science and technology.
@@
Several noteworthy findings emerge from our study. We document that even in such restricted genres as academic writing and invention abstracts, female authors tend to use more “involved” features than male authors, a pattern that holds across all scientific fields and patent subcategories. In the patent data specifically, we find that the gender of the lawyer appears to have more impact on the writing than the gender of the inventor, suggesting both the importance of attorneys in crafting patent text and the universal nature of the patterns we observe, even across individuals with very different professional training. Finally, we show that work that uses a higher number of “involved” features is more likely to be cited by female authors.
@@
Our study builds on, but also extends, a small but growing stream of research in the Science of Science on gender and language (c.f., fortunato2018, fortunato2018). Prior work in this area has yielded mixed findings regarding gender differences in scientific writing style. Some studies report notable differences between men’s and women’s writing (lerchenmueller2019; kolev2020; kim2022), while other findings suggest that differences are minimal or statistically insignificant, particularly in longer text formats (horbach2022; franco2021; lerchenmueller2019). Additional research suggests that academic writing styles among men and women are more similar than different (francis2001; rubin1992), and that gender-based variation may be more pronounced in less constrained tasks (Newman2008). However, the choice of features in much of this work has not been motivated by a linguistic framework, thereby complicating the interpretation of the results and making comparisons across studies difficult. In addition, existing research has been done within particular fields and on particular kinds of text (e.g., scientific papers), and therefore little is known about whether the patterns observed thus far will generalize across diverse fields and document types (such as patent data), which have diverse norms.
@@
Our study also adds to the linguistics literature. First, previous research on gender differences in language utilization has largely been done on informal or conversational text, typically focusing on speech patterns, and using corpora like emails (thomson2001), essays (mulac2006), social media (garimella2016), blogs (pennebaker2011). There has been far less research on gender differences in formal writing, and the work that has been done has relied on relatively small samples (argamon2003; biber1988). Second, we study a new source of text data—patented inventions—that has not been studied in previous linguistic analyses. Patents represent a more restricted genre in terms of stylistic expression than academic papers, therefore, finding stylistic differences by gender is especially notable.
2 Materials and Methods
2.1 Framework
We draw on a previously established linguistic framework, known as “Informational versus Involved Production,” outlined by biber1988, which characterizes the co-occurrence of certain linguistic features based on their communicative function. While the linguistics community has developed a number of potentially applicable frameworks, we chose Biber’s because it was created using a diverse set of corpora, including non-fiction, and was intended to capture universal features of English language; therefore, we anticipate that it would likely capture useful variation across a broad range of scientific and technological fields, encompassing potentially diverse norms of composition. We specifically chose the “informational”/“involved” dimension of Biber’s framework, as studies note that men tend to focus more on relaying factual information (“informational”), while women tend to focus on building a relationship with an audience in their communications (“involved”; argamon2003, argamon2003). Based on a corpus of e-mail communication, speech, and blogs, pennebaker2011 stipulated that “males categorize their worlds by counting, naming, and organizing the objects they confront” (informational aspect), while women “personalize their topics” (involved aspect).
Biber outlines two factors that influence the choice of individuals to use a specific linguistic mode: the production circumstances of real-time constraints (speech or conversation) versus editing possibilities (academic writing), and the primary purpose of the writer (informational versus involved). With respect to the production circumstances, given the type of corpus that we are exploring in this study (formal written text), we can hold these constant, as both papers and patents are produced in settings where editing is possible and likely (i.e., the authors are under no real-time constraints, unlike the case of a conversation). With respect to the primary purpose of the writer (informational versus involved), argamon2003 noted that formal written text is intended for a “broad unseen audience.” biber_1998 found that letters written by women and addressed to women were more involved and letters written by men and addressed to men were less involved. He also found that, overall, letters written by women are more involved than letters written by men. Even though in formal written text there is no specific addressee, and academic papers and patents are intended for a “broad unseen audience”, our goal is to understand if we are still able to observe the variation by the gender of the writers.
In addition to Biber’s framework , the specific set of “involved” and “informational” features used in our study was informed by the follow-up work of pennebaker2011 and argamon2003. We chose three distinctive features for “involved” and “informational” writing modes (Figure 2). Pronouns, questions, and non-phrasal coordination (“and” connector) are highly correlated as linguistic features and are reflective of the “involved” writing style. The use of pronouns indicates interaction between the writer and the reader—the writer assumes that the reader knows what they are referring to with a pronoun, as the reader is following the writer’s story in a text; pronouns “present things in a relational way” (argamon2003, p. 7). We consider questions to be an important involved feature as well, as they point to the presence of a hypothetical addressee in a formal written text, and are therefore a marker of interaction (biber1986; marckworthbaker1974). Non-phrasal coordination (“and”) is used “to string clauses together in a loose, logically unspecified manner, instead of integrating the information into fewer units” (biber1988, p. 106), and thus is more associated with a narrative type of communication (“involved” aspect). For “informational” writing mode indicators, we use past tense, determiners, and cardinal numbers. We use past tense as an “informational” feature, to indicate a “disconnection” with the audience, absence of interaction in the narrative. The other two “informational” features—determiners and cardinal numbers—are taken from argamon2003 who used the Balanced Winnow algorithm to identify features that are strong indicators of either male or female writing. Argamon and colleagues note that, based on their examination of texts, “the use of determiners reflects that male writers are mentioning classes of things, in contrast to female writers, who are personalizing their messages” (argamon2003, p. 12). The greater use of cardinal numbers (quantification) is also noted as a strong male indicator in both the work of pennebaker2011 and argamon2003.
2.2 Corpus
As noted previously, we draw on two different corpora: (1) abstracts of academic papers from Web of Science [^2] Web of Science: https://clarivate.com/webofsciencegroup/solutions/web-of-science/ (“WoS data”) and (2) abstracts of patents from PatentsView (“PatentsView data”) [^3] PatentsView: http://www.patentsview.org/. Both databases have been widely used in prior research in the Science of Science, and provide generally representative coverage of scientific papers and patented inventions, respectively (funk2017; birkle2020; jaffe2002; kedrick2024cp).
For the WoS data, we excluded document types that are not typically meant for the publication of scientific findings (e.g., book reviews, letters, music scores, and bibliographies). We further limited our attention to articles that included full-text abstracts, and that were written in English [^4] The WoS data do not include accurate information on the language of paper abstracts. We identified English language abstracts using a commercial machine translation API.. For the PatentsView data, we limited our attention to utility patents, which encompass roughly 90% of the patents granted by the United States Patent and Trademark Office (USPTO), and have therefore been the central focus of prior work (jaffe2002). All patents granted by the USPTO are written in English, and more than 90% (going back to 1976) include full-text abstracts.
For the purposes of our study, we further narrow our focus to single-authored papers and patents (i.e., we exclude papers and patents written by teams). Our motivation for doing so was to allow for a clearer identification of the relationship between gender and writing style, which would be more complex in the case of teams, particularly those of mixed gender. Note that in the patent data, we also narrowed our sample to include only those inventors with a single lawyer, in order to more cleanly control for gender. In addition, we require that abstracts have more than 100 words, in order to normalize frequency counts per 100 words, which is consistent with prior research in linguistics (biber1988). Additional details on the procedure used to narrow down our samples from the population of articles and patents, including the number of observations lost at each stage, are given in Figure LABEL:fig:Corpus_Size in the Supplementary Materials.
For the WoS data, we subset the corpus to years starting in 1991, as abstracts are not well represented in earlier periods (Figure 1). After matching, most of the abstracts in our final analytical sample are from the Social Sciences, followed by Life Sciences and Biomedicine. (Figure LABEL:fig:academic_papers_all_field in the Supplementary Materials shows the distribution of fields in the original Web of Science corpus.) This is consistent with prior work, as women are well represented in the social sciences and biomedical domains (ceci2014; su2015). For the PatentsView data, our analysis begins in 1976, which is the first year in which machine readable patent records are available from the USPTO. After matching, most patents come from the “Others” NBER category, which includes “Miscellaneous” patent types. (Figure LABEL:fig:Patents_full_category in the Supplementary Materials shows the distribution of categories in the original PatentsView corpus.)
2.3 Linguistic Measures
For each paper and patent in our sample, we constructed a series of measures tracking the utilization of informational and involved features. We use pronouns, questions, and non-phrasal coordination (“and” connector) as indicators of the “involved” style of writing, and past tense, determiners, and cardinal numbers as indicators of “informational” style of writing. Figure 2 gives an overview of the specific features of interest; more details about the computation of these features are included in the Table LABEL:table_features in the Supplementary Materials. [^6] We also released an open-source Python package to compute these linguistic features, available here: https://pypi.org/project/stylometer.
We define three separate measures: involved rate, informational rate, and Involved-Informational Ratio. Because our corpus includes academic writing and invention abstracts, there is already a high utilization of informational features by both female and male authors; therefore, in addition to involved and informational rates, we are also interested in seeing the relative rate at which writers use involved features in relation to informational features (Involved-Informational Ratio). We define the involved and informational rates as the sum of three involved (or informational) linguistic features divided by the total number of tokens in a given abstract, and normalized to units of 100 words (biber1988). We define the Involved-Informational Ratio as the involved rate divided by the informational rate.
$\textrm{Involved Rate}=\frac{N_{pron}+N_{and}+N_{q}}{N_{tokens}}\times 100$ (1)
$\textrm{Informational Rate}=\frac{N_{det}+N_{past}+N_{num}}{N_{tokens}}\times 100$ (2)
$\textrm{Inv.-Inf. Rate}=\frac{\textrm{Involved Rate}}{\textrm{Informational Rate}}$ (3)
Figure 1: Distribution of papers and patents by year and by field. These panels represent data used in the analytical sample. Figure LABEL:fig:academic_papers_all_field in the Supplementary Materials shows the distribution of categories in the original Web of Science corpus, and Figure LABEL:fig:Patents_full_category in the Supplementary Materials shows the distribution of categories in the original PatentsView corpus.
2.4 Citation measures
The possibility that men and women write in systematically different ways naturally raises questions about the relationship between writing styles and the reception of intellectual products. Specifically, differences in how men and women present their research in written text may help account for some of the disparities in rates of citation documented in prior work (caplar2017; potthoff2017; dion2018; dworkin2020; zurn2020).
To evaluate this possibility, we conducted a series of analyses in which we decomposed the number of citations made by future papers and patents to those in our sample, according to the gender of the citing author or inventor. We emphasize that citation is merely a proxy for (unobservable) reading preferences—authors may cite prior work without reading it; scientists and inventors probably only cite a fraction of the prior work that they do read; and citation need not correspond to a favorable reading of the cited work. Keeping those caveats in mind, for each sample paper and patent, we computed two related measures: (1) citations from papers with a female (male) first author, (2) citations from papers with a female (male) last author. For each measure, we excluded self-citations, i.e., those made by the author of the sample paper to his or her own work. In addition, to facilitate interpretation, we transform each measure into a rate per 100 citations. Subsequently, we estimated a series of regression models, wherein the four measures are our outcomes of interest, and the predictors are the gender of the author or inventor (and lawyer for patents) and the rates of informational and involved features. For papers and patents cited 0 times, we impute a rate of 0, and add an imputation indicator as a covariate. As with our prior regression, we control for the field and year of academic paper/patent.
To assess whether writing style influences citation patterns beyond content similarity, we constructed a content-based risk set of potential citers for each single-authored paper and patent. This dataset included both observed citations and up to ten counterfactual citations—papers and patents that were highly similar in content (based on SBERT embeddings) but did not cite the work. We identified potential counterfactual citers by using FAISS to retrieve the top 10 nearest neighbors (based on cosine similarity between SBERT embeddings) within the same field and publication year as the observed citers. This design enabled us to test whether involved and informational rates predict citation patterns when we control for semantic similarity. We examined gender differences in citation patterns using regression models that estimated the likelihood that a citation came from (1) a paper with a female (male) first author, or (2) a paper with a female (male) last author. The dependent variable was coded as 1 if the citation was observed and the citing paper met the gender criterion, and 0 otherwise. All models included fixed effects for field and publication year, and we clustered standard errors by paper or patent to account for variation in the number of observed and counterfactual citations. For additional details, see the Controlling for Content Similarity section in the Supplementary Materials.
@@ 3.1 Data Overview
We conduct the analysis on two datasets: abstracts of academic papers from Web of Science [^7] Web of Science: https://clarivate.com/webofsciencegroup/solutions/web-of-science/ and abstracts of patents from PatentsView [^8] PatentsView: http://www.patentsview.org/. For the purposes of our study, we narrow our focus to single-authored papers and patents (i.e., we exclude papers and patents written by teams). In the patent data, we also narrow our sample to include only those inventors with a single lawyer, in order to more cleanly control for the gender of the lawyer, as lawyers are frequently involved in the writing and editing of patent text. In total, we use 256,260 publications and 32,106 patents in the analysis. Additional details on the dataset construction can be found in the Materials and Methods section.
For each paper and patent in our sample, we construct a series of measures tracking the utilization of informational features (using past tense, determiners, and cardinal numbers as indicators) and involved features (using pronouns, questions, and non-phrasal coordination (“and”) as indicators). Figure 2 gives an overview of the specific features of interest as well as more details about the computation of these features. [^9] We also released an open-source Python package to compute these linguistic features, available here: https://pypi.org/project/stylometer. We define three separate measures: involved rate, informational rate, and Involved-Informational Ratio. Because our corpus includes academic writing and invention abstracts, there is already a high utilization of informational features; therefore, in addition to involved and informational rates, we are also interested in seeing the relative rate at which writers use involved features in relation to informational features (Involved-Informational Ratio). More information on the developed measures is included in the Materials and Methods section.
@@ 3.2 Writing Examples
Although both abstracts address similar topics, they are presented in dramatically different ways. Abstract (2), written by the male author, is striking for its heavy utilization of cardinal numbers (e.g., “\$100 million”, “5.5 years”, “above 336.69”), which are mostly avoided by the female author. Abstract (1), by contrast, stands out for its greater utilization of “connector words” (e.g., “and”) and questions, the latter of which are not used at all in Abstract (2). The female abstract reads a lot like a story, building a relationship with the reader through the use of pronouns, present tense, asking questions, and by stringing the clauses together with “and” in a loose, storytelling-like manner; the male abstract is focused on relaying the factual information, via the use of numbers and determiners, and past tense. Figure 3 shows the full text of both abstracts, zoomed out, with informational and involved features highlighted in different colors (blue for informational, orange for involved).